# Decentralized Proactive Model Offloading and Resource Allocation for Split and Federated Learning

### v1

@@ 3.2 The latency of DNN model training adopting split federated learning
(1) Starting Phase. The starting phase mainly includes device-side model distribution. At the beginning of training round $t$, the latest device-side sub-model $\mathbf{w}_{d}^{n}(t)$ is distributed to the IoT device $I_{n}$ by the edge server. The device-side sub-model’s size (in bits) for each IoT device $I_{n}$ is $\psi_{d,n}^{m}(l_{n})$, depending on cut layer $l_{n}$. The downlink transmission rate $r_{d,n}^{DL}$ from the edge server to IoT device $I_{n}$ can be calculated by
@@
where $\mu_{d,n}^{DL}$ is the time fraction allocated to the IoT device $I_{n}$ on the downlink channel. $W^{DL}$ is the radio spectrum bandwidth for the edge server’s uplink channel. $P_{s}$ is the edge server’s transmission power. $h_{d}^{n}$ is the channel gain between the edge server and IoT device $I_{n}$. $N_{0}$ is the thermal noise spectrum density. Hence, the distribution latency $\tau_{d,n}^{m,DL}$ of device-side sub-model $\mathbf{w}_{d}^{n}$ for IoT device $I_{n}$ can be calculated by
@@
(2) Intermediate Phase. After receiving its latest device-side model, each IoT device $I_{n}$ trains its DNN model $\mathbf{w}_{n}$ for $\Upsilon$ epochs in training round $t$, indexed by $\upsilon\in\{1,...,\upsilon,...,\Upsilon\}$. At the first epoch in training round $t$, we have $\mathbf{w}_{d}^{n}(t,1)\leftarrow\mathbf{w}_{d}^{n}(t),\forall I_{n}\in\mathcal{I}$. For each IoT device $I_{n}$, its model training process at each epoch $\upsilon$ in training round $t$ consist of device-side model execution, smashed data transmission, service-side model execution, service-side model update, smashed data’s gradient transmission and device-side model update six parts. The latency of each part can be analyzed as follows.
• Device-side model execution latency. It represents the time taken by the device-side model $\mathbf{w}_{d}^{n}$ of IoT device $I_{n}$ performing forward propagation on a mini-batch data samples. At an epoch $\upsilon$ in training round $t$, each IoT device $I_{n}$ randomly draws a mini-batch of data samples $\mathcal{B}_{d}^{n}$ from its local dataset $\mathcal{D}_{d}^{n}$ and feed them into device-side model to perform forward propagation. Here, the mini-batch size for IoT device $I_{n}$ can be denoted by $B_{d}^{n}=|\mathcal{B}_{d}^{n}|$. Hence, the total batches $b_{d}^{n}$ of IoT device $I_{n}$ for an epoch is calculated by $b_{d}^{n}=D_{d}^{n}/B_{d}^{n}$. Let $\phi_{s,n}^{f,e}(l_{n})$ denote the computation workload (in FLOPs) for the device-side model of IoT device $I_{n}$ performing forward propagation on a single data sample. Hence, the latency $\tau_{d,n}^{f,e}$ for the device-side model processing a mini-batch of data samples can be calculated by
@@
• Smashed data transmission latency. It represents the time taken by IoT device $I_{n}$ transmitting the smashed data for a mini-batch of data samples to the edge server. After the device-side model performing forward propagation on a mini-batch of data samples, the smashed data of these data samples need to be transmitted to the edge server using the allocated bandwidth resource. Let $\psi_{d,n}^{s,UL}(l_{n})$ be the smashed data size of one data sample at the cut layer $l_{n}$. Hence, the smashed data size in bits for a mini-batch $B_{d}^{n}$ of data samples can be denoted by $B_{d}^{n}\psi_{d,n}^{s,UL}(l_{n})$. The number of subcarriers allocated to IoT device $I_{n}$ can be denoted by $\mu_{d}^{n}$. The uplink transmission rate $r_{d,n}^{UL}$ from IoT device $I_{n}$ to edge server can be calculated by
@@
where $W^{UL}$ is the radio spectrum bandwidth for the edge server’s downlink channel. $P_{d}^{n}$ denotes the transmission power for IoT device $I_{n}$. $h_{d}^{n}$ denotes the channel gain between IoT device $I_{n}$ and edge server. Hence, the smash data transmission latency is given by
@@
• Service-side model execution latency. It represents the time taken by the server-side model $\mathbf{w}_{s}^{n}$ of IoT device $I_{n}$ performing forward propagation on a mini-batch data samples. Let $\phi_{s,n}^{f,e}(l_{n})$ denote the computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of IoT device $I_{n}$ performing forward propagation on a single data sample. Since the smash data for a mini-batch of data samples are fed for training the server-side model, the overall computation workload is $B_{d}^{n}\phi_{s,n}^{f,e}(l_{n})$. Hence, the service-side model execution latency can be calculated by
@@
where $f_{s}$ is the computation capacity of the edge server. $\theta_{d}^{n}$ denotes the edge server’s computation capacity fraction allocated to IoT device $I_{n}$;
• Service-side model update latency. It represents the time taken by the server-side model $\mathbf{w}_{s}^{n}$ of IoT device $I_{n}$ performing backward propagation on a mini-batch data samples. Let $\phi_{s,n}^{b,e}(l_{n})$ denote the computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of IoT device $I_{n}$ performing backward propagation on a single data sample. Hence, the overall computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of IoT device $I_{n}$ performing backward propagation on a mini-batch of data samples is $B_{d}^{n}\phi_{s,n}^{b,e}(l_{n})$. The service-side model update latency can be calculated by
@@
• Smashed data’s gradient transmission latency. It represents the time taken by the edge server transmitting the smashed data’s gradients for a mini-batch of data samples to IoT device $I_{n}$. After service-side model of IoT device $I_{n}$ are updated, its smashed data’s gradients are sent back to IoT device $I_{n}$ using the allocated radio spectrum. Let $\psi_{s,n}^{g,DL}(l_{n})$ denote the data size of smashed data’s gradient for a single data sample. Hence, the smash data’s gradient transmission latency for a mini-batch data samples can be calculated by
@@
• Device-side model update latency. It refers to the time taken by the device-side model $\mathbf{w}_{d}^{n}$ of IoT device $I_{n}$ performing back propagation on a mini-batch data samples. Let $\phi_{d,n}^{b,e}(l_{n})$ represent the computation workload for the device-side model $\mathbf{w}_{d}^{n}$ of IoT device $I_{n}$ performing backward propagation on a single data sample. Hence, the device-side model update latency can be calculated by
@@
Based on the above analysis, the overall latency of IoT device $I_{n}$ at epoch $\upsilon$ in training round $t$ can be calculated by
@@
(3) End Phase. Device-side model transmission latency and model aggregation latency. After each IoT device $I_{n}$ training its DNN model $\mathbf{w}_{n}$ consisting of device-side model $\mathbf{w}_{d}^{n}$ and service-side model $\mathbf{w}_{s}^{n}$ for $\Upsilon$ epochs, the latest device-side model is transmitted from each IoT device $I_{n}$ to the edge server, and the corresponding transmission latency for each IoT device $I_{n}$ can be calculated by
@@
The model parameters are aggregated on the edge server by the FedAvg algorithm. Due to its low computational complexity, the model aggregation latency is very small and negligible.
(4) Overall training latency. With the results of three phases in (2), (10) and (11), the overall training latency of IoT device $I_{n}$ in one training round $t$ can be calculated by
$\tau_{n}^{t}=\tau_{s,n}^{m,DL}+\Upsilon\tau_{n}^{m,e}+\tau_{n,s}^{m,UL},\forall I_{n}\in\mathcal{I}$ (12)
which depends on the cut layer $l_{n}$ of DNN model on IoT device $I_{n}$, the time fraction $\mu_{d,n}^{DL}$ allocated to IoT device $I_{n}$ on the downlink channel, the time fraction $\mu_{d,n}^{UL}$ allocated to IoT device $I_{n}$ on the uplink channel, and the edge server’s computation capacity’s fraction $\theta_{d}^{n}$ allocated to IoT device $I_{n}$. Specifically, the cut layer not only affects the computation workload distribution and the transmission data amount between the IoT device and the edge server, but also affects the data leakage risk rate. Different cut layer have different amounts of transmission data. A shallow cut layer means light computational workload on IoT devices, but high data leakage risk. A deep cut layer means heavy computation workloads on IoT devices, but low data leakage risk. We further quantify the data leakage risk in next subsection.
3.3 The data leakage risk of DNN model training adopting split federated learning
Although offloading model from resource-constrained IoT devices to the edge server can reduce the training latency and improve training efficiency, the raw data can be partially reconstructed from the gradient information [16, 15], thereby incurring the data leakage risk. The recovery of data samples from gradient information is described in detail as follows:
(1) Service-side model’s gradient of the original data samples. In the first epoch of training round $t$, the edge server owns the device-side model $\mathbf{w}_{d}^{n}(t,1)$ of each IoT device $I_{n}$. Let $\mathbf{Z}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ denote the aggregated input of a mini-batch of data samples in IoT device $I_{n}$. Each IoT device $I_{n}$ executes its device-side model with the raw data samples, and obtains smashed data $\mathbf{S}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times P}$, i.e.,
@@ 3.3 The data leakage risk of DNN model training adopting split federated learning
where $f(\mathbf{Z}_{d}^{n};\mathbf{w}_{d}^{n})$ represents the mapping function between a mini-batch of data samples $\mathbf{Z}_{d}^{n}$ and their smash data $\mathbf{S}_{d}^{n}$ given device-side sub-model parameter $\mathbf{w}_{d}^{n}$. $P$ denotes the dimension of smashed data for one data sample. Device-side model $\mathbf{w}_{d}^{n}$ of each IoT device $I_{n}$ performs forward propagation on the aggregated input $\mathbf{Z}_{d}^{n}(t,1)$ of a mini-batch of data samples and transmits its smashed data $\mathbf{S}_{d}^{n}(t,1)$ to the edge server. The edge server receives the smashed data $\mathbf{S}_{d}^{n}(t,1)$ of each IoT device $I_{n}$ and feed it into the server-side model $\mathbf{w}_{s}^{n}(t,1)$. As such, the predicted result from the service-side model is given by
@@
According to the loss function between the predicted results $\mathbf{\hat{y}}_{d}^{n}$ and the corresponding ground-truth labels $\mathbf{y}_{d}^{n}$, the gradient of the service-side model can be calculated and denoted by $\nabla L(\mathbf{w}_{s}^{n}(t,1))$.
(2) Service-side model’s gradient of the recovered data samples. Let $\mathbf{Z}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ denote the mini-batch of recovered data samples in IoT device $I_{n}$. The edge server executes the device-side model of IoT device $I_{n}$ with the mini-batch of recovered data samples, and obtains corresponding smashed data $\mathbf{S}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times P}$
@@
The smashed data $\mathbf{S}_{d,n}^{{}^{\prime}}(t,1)$ is fed into the server-side model $\mathbf{w}_{s}^{n}(t,1)$ of IoT device $I_{n}$, and it outputs the predicted results $\hat{\mathbf{y}}_{d,n}^{{}^{\prime}}(t,1)$,
@@
According to the loss function between the predicted results $\mathbf{\hat{y}}_{d,n}^{{}^{\prime}}$ and the corresponding ground-truth labels $\mathbf{y}_{d}^{n}$, the service-side model’s gradient of the recovered data samples can be calculated and denoted by $\nabla L^{{}^{\prime}}(\mathbf{w}_{s}^{n}(t,1))$.
(3) Recovering data samples from their gradients. In the first epoch of each training round $t$, the edge server owns the device-side model $\mathbf{w}_{d}^{n}(t,1)$ and the service-side model $\mathbf{w}_{s}^{n}(t,1)$. Thus, the edge server try to recovery data samples by optimizing an euclidean matching term between $\nabla L(\mathbf{w}_{s}^{n}(t,1))$ and $\nabla L^{{}^{\prime}}(\mathbf{w}_{s}^{n}(t,1))$ [16, 15]. The optimization objective function
@@
According to the above analysis, we define the data leakage risk $P_{d}^{n}(l_{n})$ of cut layer $l_{n}$ for each IoT device $I_{n}$ as the cosine similarity of the original data samples $\mathbf{Z}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ and the recovered data samples $\mathbf{Z}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$,
@@
3.4 Regression-based modeling methodology
As shown in Subsection 3.2, in the overall training latency model, some functions including the forward propagation workload function with respect to the cut layer, the backward propagation workload function with respect to the cut layer, the smashed data size function with the respect to the cut layer, the smashed data’s gradient size with the respect to the cut layer, the device-side model’s data size with respect to the cut layer, cannot be expressed in an analytic form. This is because that these functions vary with different DNN mdoels. For example, the specific coefficients in the above these functions are different for Restnet 18 and Restnet 34 two models.
To address the above challenge, referring to the existing works [23, 11], we adopt a data-driven methodology. We first empirically measure the forward/backward propagation workloads, smashed data/smashed data’s gradient sizes, and device-side model’s data size with respect to different cut layers for Restnet 18 and Restnet 34 two models. Then we employ a variety of classic regression models to profile the above relationships. We find that it is optimal to adopt Quadratic Polynomial Regression (QPR) regression models to profile the functions between the forward/backward propagation workloads, the device-side model size and the cut layer. Also, it is optimal to adopt Reciprocal Regression (RR) to profile function between the smashed data/smashed data’s gradient size and the cut layer. The developed regression models for Restnet 18 and Restnet 34 two models are shown in Table I. The root mean square error (RMSE) is applied for calculating the average model-prediction error in the units of the variable of interest [24].
DNN Models
Functions Proposed models RMSE
Restnet 18
$\psi^{m}(l_{n})$ $6.5280.9746x^{2}-5.58x+$ 6.528 3.235

### v2

@@ III-B The latency of DNN model training adopting split federated learning
(1) Starting Phase. The starting phase mainly includes device-side model distribution. At the beginning of training round $t$, the latest device-side sub-model $\mathbf{w}_{d}^{n}(t)$ is distributed to the end device $I_{n}$ by the edge server. The device-side sub-model’s size (in bits) for each end device $I_{n}$ is $\psi_{d,n}^{m}(l_{n})$, depending on cut layer $l_{n}$. The downlink transmission rate $r_{d,n}^{DL}$ from the edge server to end device $I_{n}$ can be calculated by
@@
where $\mu_{d,n}^{DL}$ is the time fraction allocated to the end device $I_{n}$ on the downlink channel. $W^{DL}$ is the radio spectrum bandwidth for the edge server’s uplink channel. $P_{s}$ is the edge server’s transmission power. $h_{d}^{n}$ is the channel gain between the edge server and end device $I_{n}$. $N_{0}$ is the thermal noise spectrum density. Hence, the distribution latency $\tau_{d,n}^{m,DL}$ of device-side sub-model $\mathbf{w}_{d}^{n}$ for end device $I_{n}$ can be calculated by
@@
(2) Intermediate Phase. After receiving its latest device-side model, each end device $I_{n}$ trains its DNN model $\mathbf{w}_{n}$ for $\Upsilon$ epochs in training round $t$, indexed by $\upsilon\in\{1,\cdots,\upsilon,\cdots,\Upsilon\}$. At the first epoch in training round $t$, we have $\mathbf{w}_{d}^{n}(t,1)\leftarrow\mathbf{w}_{d}^{n}(t),\forall I_{n}\in\mathcal{I}$. For each end device $I_{n}$, its model training process at each epoch $\upsilon$ in training round $t$ consists of device-side model execution, smashed data transmission, server-side model execution, server-side model update, smashed data’s gradient transmission and device-side model update six parts. The latency of each part can be analyzed as follows.
• Device-side model execution latency. It represents the time taken by the device-side model $\mathbf{w}_{d}^{n}$ of end device $I_{n}$ performing forward propagation on a mini-batch data samples. At an epoch $\upsilon$ in training round $t$, each end device $I_{n}$ randomly draws a mini-batch of data samples $\mathcal{B}_{d}^{n}$ from its local dataset $\mathcal{D}_{d}^{n}$ and feed them into device-side model to perform forward propagation. Here, the mini-batch size for end device $I_{n}$ can be denoted by $B_{d}^{n}=|\mathcal{B}_{d}^{n}|$. Hence, the total batches $b_{d}^{n}$ of end device $I_{n}$ for an epoch is calculated by $b_{d}^{n}=D_{d}^{n}/B_{d}^{n}$. Let $\phi_{s,n}^{f,e}(l_{n})$ denote the computation workload (in FLOPs) for the device-side model of end device $I_{n}$ performing forward propagation on a single data sample. Hence, the latency $\tau_{d,n}^{f,e}$ for the device-side model processing a mini-batch of data samples can be calculated by
@@
• Smashed data transmission latency. It represents the time taken by end device $I_{n}$ transmitting the smashed data for a mini-batch of data samples to the edge server. After the device-side model performing forward propagation on a mini-batch of data samples, the smashed data of these data samples need to be transmitted to the edge server using the allocated bandwidth resource. Let $\psi_{d,n}^{s,UL}(l_{n})$ be the smashed data size of one data sample at the cut layer $l_{n}$. Hence, the smashed data size in bits for a mini-batch $B_{d}^{n}$ of data samples can be denoted by $B_{d}^{n}\psi_{d,n}^{s,UL}(l_{n})$. The number of subcarriers allocated to end device $I_{n}$ can be denoted by $\mu_{d}^{n}$. The uplink transmission rate $r_{d,n}^{UL}$ from end device $I_{n}$ to edge server can be calculated by
@@
where $W^{UL}$ is the radio spectrum bandwidth for the edge server’s downlink channel. $P_{d}^{n}$ denotes the transmission power for end device $I_{n}$. $h_{d}^{n}$ denotes the channel gain between end device $I_{n}$ and edge server. Hence, the smash data transmission latency is given by
@@
• Server-side model execution latency. It represents the time taken by the server-side model $\mathbf{w}_{s}^{n}$ of end device $I_{n}$ performing forward propagation on a mini-batch data samples. Let $\phi_{s,n}^{f,e}(l_{n})$ denote the computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of end device $I_{n}$ performing forward propagation on a single data sample. Since the smash data for a mini-batch of data samples are fed for training the server-side model, the overall computation workload is $B_{d}^{n}\phi_{s,n}^{f,e}(l_{n})$. Hence, the server-side model execution latency can be calculated by
@@
where $f_{s}$ is the computation capacity of the edge server. $\theta_{d}^{n}$ denotes the edge server’s computation capacity fraction allocated to end device $I_{n}$.
• Server-side model update latency. It represents the time taken by the server-side model $\mathbf{w}_{s}^{n}$ of end device $I_{n}$ performing backward propagation on a mini-batch data samples. Let $\phi_{s,n}^{b,e}(l_{n})$ denote the computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of end device $I_{n}$ performing backward propagation on a single data sample. Hence, the overall computation workload for the server-side model $\mathbf{w}_{s}^{n}$ of end device $I_{n}$ performing backward propagation on a mini-batch of data samples is $B_{d}^{n}\phi_{s,n}^{b,e}(l_{n})$. The server-side model update latency can be calculated by
@@
• Smashed data’s gradient transmission latency. It represents the time taken by the edge server transmitting the smashed data’s gradients for a mini-batch of data samples to end device $I_{n}$. After server-side model of end device $I_{n}$ is updated, its smashed data’s gradients are sent back to end device $I_{n}$ using the allocated radio spectrum. Let $\psi_{s,n}^{g,DL}(l_{n})$ denote the data size of smashed data’s gradient for a single data sample. Hence, the smash data’s gradient transmission latency for a mini-batch data samples can be calculated by
@@
• Device-side model update latency. It refers to the time taken by the device-side model $\mathbf{w}_{d}^{n}$ of end device $I_{n}$ performing backpropagation on a mini-batch data samples. Let $\phi_{d,n}^{b,e}(l_{n})$ represent the computation workload for the device-side model $\mathbf{w}_{d}^{n}$ of end device $I_{n}$ performing backward propagation on a single data sample. Hence, the device-side model update latency can be calculated by
@@
Based on the above analysis, the overall latency of end device $I_{n}$ at epoch $\upsilon$ in training round $t$ can be calculated by
@@
(3) End Phase. The end phase is to aggregate the latest device-side models to the edge server to obtain the updated global DNN model. Therefore, the latency for the end phase is mainly consisting of the device-side model transmission latency and the model parameters aggregation latency on the edge server. After each end device $I_{n}$ training its DNN model $\mathbf{w}_{n}$ consisting of device-side model $\mathbf{w}_{d}^{n}$ and server-side model $\mathbf{w}_{s}^{n}$ for $\Upsilon$ epochs, the latest device-side model is transmitted from each end device $I_{n}$ to the edge server, and the corresponding transmission latency $\tau_{n,s}^{m,UL}$ for each end device $I_{n}$ can be calculated by
@@
where $\psi_{d,n}^{m}(l_{n})$ is the device-side sub-model size (in bits) for each end device $I_{n}$. $r_{d,n}^{UL}$ is the uplink transmission rate from the end device to the edge serve. The model parameters are aggregated on the edge server by the FedAvg algorithm. Due to its low computational complexity, the model aggregation latency is very small and negligible.
(4) Overall training latency. With the results of three phases in (2), (10) and (11), the overall training latency of end device $I_{n}$ in one training round $t$ can be calculated by
$\tau_{n}^{t}(l_{n},\mu_{d,n}^{DL},\mu_{d,n}^{UL},\theta_{d}^{n})=\tau_{s,n}^{m,DL}+\Upsilon\tau_{n}^{m,e}+\tau_{n,s}^{m,UL},\forall I_{n}\in\mathcal{I}$ (12)
which depends on the cut layer $l_{n}$ of DNN model on end device $I_{n}$, the time fraction $\mu_{d,n}^{DL}$ allocated to end device $I_{n}$ on the downlink channel, the time fraction $\mu_{d,n}^{UL}$ allocated to end device $I_{n}$ on the uplink channel, and the edge server’s computation capacity’s fraction $\theta_{d}^{n}$ allocated to end device $I_{n}$. Specifically, the cut layer affects not only the computation workload distribution and the transmission data amount between the end device and the edge server but also the data leakage risk rate. Different cut layers have different amounts of transmission data. A shallow cut layer means light computational workload on end devices, but high data leakage risk. A deep cut layer means heavy computation workloads on end devices, but low data leakage risk. We further quantify the data leakage risk in next subsection.
III-C The data leakage risk of DNN model training adopting split federated learning
Although offloading model from resource-constrained end devices to the edge server can reduce the training latency and improve training efficiency, the raw data can be partially reconstructed from the gradient information [16, 15], thereby incurring the data leakage risk. The recovery of data samples from gradient information is described in detail as follows:
(1) Server-side model’s gradient of the original data samples. In the first epoch of training round $t$, the edge server owns the device-side model $\mathbf{w}_{d}^{n}(t,1)$ of each end device $I_{n}$. Let $\mathbf{Z}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ denote the aggregated input of a mini-batch of data samples in end device $I_{n}$. Each end device $I_{n}$ executes its device-side model with the raw data samples, and obtains smashed data $\mathbf{S}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times P}$, i.e.,
@@ III-C The data leakage risk of DNN model training adopting split federated learning
where $f(\mathbf{Z}_{d}^{n};\mathbf{w}_{d}^{n})$ represents the mapping function between a mini-batch of data samples $\mathbf{Z}_{d}^{n}$ and their smash data $\mathbf{S}_{d}^{n}$ given device-side sub-model parameter $\mathbf{w}_{d}^{n}$. $P$ denotes the dimension of smashed data for one data sample. Device-side model $\mathbf{w}_{d}^{n}$ of each end device $I_{n}$ performs forward propagation on the aggregated input $\mathbf{Z}_{d}^{n}(t,1)$ of a mini-batch of data samples and transmits its smashed data $\mathbf{S}_{d}^{n}(t,1)$ to the edge server. The edge server receives the smashed data $\mathbf{S}_{d}^{n}(t,1)$ of each end device $I_{n}$ and feeds it into the server-side model $\mathbf{w}_{s}^{n}(t,1)$. As such, the predicted result from the server-side model is given by
@@
According to the loss function between the predicted results $\mathbf{\hat{y}}_{d}^{n}$ and the corresponding ground-truth labels $\mathbf{y}_{d}^{n}$, the gradient of the server-side model can be calculated and denoted by $\nabla L(\mathbf{w}_{s}^{n}(t,1))$.
(2) Server-side model’s gradient of the recovered data samples. Let $\mathbf{Z}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ denote the mini-batch of recovered data samples in end device $I_{n}$. The edge server executes the device-side model of end device $I_{n}$ with the mini-batch of recovered data samples, and obtains corresponding smashed data $\mathbf{S}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times P}$
@@
The smashed data $\mathbf{S}_{d,n}^{{}^{\prime}}(t,1)$ is fed into the server-side model $\mathbf{w}_{s}^{n}(t,1)$ of end device $I_{n}$, and it outputs the predicted results $\hat{\mathbf{y}}_{d,n}^{{}^{\prime}}(t,1)$,
@@
According to the loss function between the predicted results $\mathbf{\hat{y}}_{d,n}^{{}^{\prime}}$ and the corresponding ground-truth labels $\mathbf{y}_{d}^{n}$, the server-side model’s gradient of the recovered data samples can be calculated and denoted by $\nabla L^{{}^{\prime}}(\mathbf{w}_{s}^{n}(t,1))$.
(3) Recovering data samples from their gradients. In the first epoch of each training round $t$, the edge server owns the device-side model $\mathbf{w}_{d}^{n}(t,1)$ and the server-side model $\mathbf{w}_{s}^{n}(t,1)$. Thus, the edge server trys to recover data samples by optimizing an euclidean matching term between $\nabla L(\mathbf{w}_{s}^{n}(t,1))$ and $\nabla L^{{}^{\prime}}(\mathbf{w}_{s}^{n}(t,1))$ [16, 15]. The optimization objective function
@@
According to the above analysis, we define the data leakage risk $P_{d}^{n}(l_{n})$ of cut layer $l_{n}$ for each end device $I_{n}$ as the cosine similarity of the original data samples $\mathbf{Z}_{d}^{n}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$ and the recovered data samples $\mathbf{Z}_{d,n}^{{}^{\prime}}(t,1)\in\mathbb{R}^{B_{d}^{n}\times Q}$,
@@
III-D Regression-based modeling methodology
As shown in Subsection 3.2, in the overall training latency model, some functions including the forward propagation workload function with respect to the cut layer, the backward propagation workload function with respect to the cut layer, the smashed data size function with respect to the cut layer, the smashed data’s gradient’s size with the respect to the cut layer, the device-side model’s data size with respect to the cut layer, cannot be expressed in an analytic form. This is because these functions vary with different DNN models. For example, the specific coefficients in the above functions are different for ResNet 18 and ResNet 34 two models.
To address the above challenges, referring to the existing works [23, 11], we adopt a data-driven methodology. We first empirically measure the forward/backward propagation workloads, smashed data/smashed data’s gradient sizes, and device-side model’s data size with respect to different cut layers for ResNet 18 and ResNet 34 two models. Then we employ a variety of classic regression models to profile the above relationships. We find that it is optimal to adopt Quadratic Polynomial Regression (QPR) regression models to profile the functions between the forward/backward propagation workloads, the device-side model size and the cut layer. Also, it is optimal to adopt Reciprocal Regression (RR) to profile the function between the smashed data/smashed data’s gradient size and the cut layer. The developed regression models for ResNet 18 and ResNet 34 two models are shown in Table II.
The root means square error (RMSE) is applied for calculating the average model-prediction error in the units of the variable of interest [24].
TABLE II: The Proposed Regression-based Models
DNN Models Functions Proposed models RMSE
ResNet 18 $\psi^{m}(l_{n})$ $6.5280.9746x^{2}-5.58x+$ 6.528 3.235